If c-a=mb then: cot(C-A)/2=(1+mcosC)/ (msinC). How?
1
Expert's answer
2017-02-21T09:38:36-0500
For example :

C = Pi/2
A = 0
m = 2; b=1
cot(Pi/2) = 0, but [1+2*cosPi/2 ] /( 2*Sin[Pi/2] )= 1/2, 1/2 not equal 0

If c-a=mb then: cot(C-A)/2=(1+mcosC)/ (msinC) - false